Determine the magnitude of the complex number z = 3 – j * 4 using the...

  1. Determine the magnitude of the complex number z = 3 – j * 4 using the complex conjugate notation and determine its phase angle.

  1. Convert the complex number z = 2.5 * e ^ (j * 60o) to the rectangular notation using the relationship between real and imaginary parts and magnitude and phase angle.

Please answer in MATLAB
Answer #1

z = 3-4j

mag= abs(z) %magnitude

Ph = angle(z) %phase angle

z2 = 2.5*(cosd(60)+1j*sind(60)) %to rectangular form